40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 26 03:57:06 2009, rob, Update, Locking, bad 
    Reply  Tue Oct 27 18:19:14 2009, rob, Update, Locking, bad MCF_issue.pngno_MCF_issue.png
       Reply  Thu Oct 29 21:51:07 2009, rob, Update, Locking, bad itsback.png
Message ID: 2162     Entry time: Thu Oct 29 21:51:07 2009     In reply to: 2152
Author: rob 
Type: Update 
Category: Locking 
Subject: bad 

Quote:

Quote:

Lock acquisition has gone bad tonight. 

The initial stage works fine, up through handing off control of CARM to MCL.  However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal.  Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost. 

 This problem has disappeared.  I don't know what it was. 

The first plot shows one of the symptoms.  The second plot is a similar section taken from a more normal acquisition sequence the night before.

All is not perfect, however, as now the handoff to RF CARM is not working.

 

The problem has returned.  I still don't know what it is, but it's making me angry. 

Attachment 1: itsback.png  13 kB  | Show | Show all
ELOG V3.1.3-